Find the coodinates of the points one-third and two-thirds of the way from the point (1,2) to the point (7,11) in the coodinate plane.
Answer by venugopalramana(3286)

Find the coodinates of the points one-third and two-thirds of the wa from the point (1,2) to the point (7,11) in the coodinate plane.
formula is
x=(mx2+nx1)/(m+n)..............y=(my2+ny1)/(m+n)...where
P(x,y) divides AB....A(x1,y1)....and ...b(x2,y2)...inthe ration m:n
we m/n=1/2
x=(1*7+2*1)/(1+2)=3
y=(1*11+2*2)/(1+2)=5
P is (3,5)